a) Convert 145 to binary,

b) Convert 11001100 to decimal,

c) convert 01101011 to hex,

d) convert C7 to binary and then binary to decimal
Answer #1

`Hey,

Note: If you have any queries related the answer please do comment. I would be very happy to resolve all your queries.

a)


145 / 2 = 72 with 1 remainder
72 / 2 = 36 with 0 remainder
36 / 2 = 18 with 0 remainder
18 / 2 = 9 with 0 remainder
9 / 2 = 4 with 1 remainder
4 / 2 = 2 with 0 remainder
2 / 2 = 1 with 0 remainder
1 / 2 = 0 with 1 remainder

So,

10010001

b)

(11001100)2 = (204)10

STEP BY STEP SOLUTION

Step 1: Write down the binary number:

11001100

Step 2: Multiply each digit of the binary number by the corresponding power of two:

1x27 + 1x26 + 0x25 + 0x24 + 1x23 + 1x22 + 0x21 + 0x20

Step 3: Solve the powers:

1x128 + 1x64 + 0x32 + 0x16 + 1x8 + 1x4 + 0x2 + 0x1 = 128 + 64 + 0 + 0 + 8 + 4 + 0 + 0

Step 4: Add up the numbers written above:

128 + 64 + 0 + 0 + 8 + 4 + 0 + 0 = 204.

So, 204 is the decimal equivalent of the binary number 11001100.

Note: Brother According to Chegg's policy we are only allowed to answer first 2 part if there are many. So, I request you to post other part as separate posts

Kindly revert for any queries

Thanks.